# DynamicTree: Interactive Real Tree Animation via Sparse Voxel Spectrum

### v1

@@ Abstract
Generating dynamic and interactive 3D objects, such as trees, has wide applications in virtual reality, games, and world simulation. Nevertheless, existing methods still face various challenges in generating realistic 4D motion for complex real trees. In this paper, we propose DynamicTree, the first framework that can generate long-term, interactive animation of 3D Gaussian Splatting trees. Unlike prior optimization-based methods, our approach generates dynamics in a fast feed-forward manner. The key success of our approach is the use of a compact sparse voxel spectrum to represent the tree movement. Given a 3D tree from Gaussian Splatting reconstruction, our pipeline first generates mesh motion using the sparse voxel spectrum and then binds Gaussians to deform the mesh. Additionally, the proposed sparse voxel spectrum can also serve as a basis for fast modal analysis under external forces, allowing real-time interactive responses. To train our model, we also introduce 4DTree, the first large-scale synthetic 4D tree dataset containing 8,786 animated tree meshes with semantic labels and 100-frame motion sequences. Extensive experiments demonstrate that our method achieves realistic and responsive tree animations, significantly outperforming existing approaches in both visual quality and computational efficiency. https://dynamictree-dev.github.io/DynamicTree.github.io/
@@ 1 Introduction
With recent advances of neural radiance fields (NeRF) Mildenhall et al. (2021) and 3D Gaussian Splatting (3DGS) Kerbl et al. (2023), high-quality reconstruction and real-time rendering become feasible. Driven by these, there is high demand to make static reconstructions interactable, for immersive experiences like 3D games, movies, and virtual reality Jiang et al. (2024); Franke et al. (2025); Schieber et al. (2025). As a vital component of natural landscapes, tree animation can significantly enrich immersive digital experiences. For example, when viewing a reconstructed backyard on a VR headset, if trees can sway gently in the wind or respond to dragging interaction, it would significantly enhance immersion.
@@
However, animating a realistic 3D tree remains challenging. Traditional tree animation methods Quigley et al. (2017); Pirk et al. (2017; 2014) typically construct physical tree models and then perform dynamic simulations. Although such approaches can generate realistic motion dynamics, they are either confined to synthetic tree models or require labor-intensive mesh refinements for high-quality rendering. Consequently, they are ill-suited for animating reconstructed 3DGS representations to faithfully reproduce real-world experiences.
Figure 1: DynamicTree achieves long-term, realistic animation and dynamic interaction with real reconstructed 3DGS trees, which first generates mesh motion via a compact sparse voxel spectrum representation and then deforms the surface-bound Gaussian primitives. We visualize the slice of the generated motion at the orange scanline along the time dimension.
@@
To animate 3DGS trees, existing methods can be broadly categorized into 4D generation and physics-based simulation. 4D generation approaches first create a static 3DGS model from text or images Bahmani et al. (2024); Singer et al. (2023); Ling et al. (2024); Zheng et al. (2024); Xu et al. (2024); Miao et al. (2024); Yin et al. (2023), then use video diffusion models (VDMs) to optimize the 4D representation. These methods, however, involve costly per-scene optimization and often yield short sequences with poor geometric consistency. In contrast, physics-based approaches, such as PhysGaussian Xie et al. (2024), couple 3DGS with physical simulation engines like Material Point Method (MPM) Jiang et al. (2015), achieving better 3D consistency and longer motion sequences. Nonetheless, they rely on simplified assumptions (e.g., uniform material), which reduce realism and are computationally expensive, making them unsuitable for real-time applications.
@@
In this work, we introduce DynamicTree, a novel 4D generative framework for animating 3DGS trees with real-time interactions. Unlike VDM-based 4D generation methods, DynamicTree directly learns tree motion priors in 3D space, avoiding geometric inconsistency. Moreover, compared to computationally expensive physics-based optimization, our generative model directly synthesizes 3D tree motion in a feed-forward manner, achieving more than a hundred times acceleration. To train the model, we construct the first large-scale 4D tree motion dataset with $8,786$ animated meshes, semantic labels, and 100-frame sequences, generated via hierarchical branching simulation Weber & Penn (1995).
@@
Our framework addresses these challenges with a two-stage pipeline. We first generate mesh motion and then bind Gaussians to the deforming mesh, allowing full 3DGS deformation while only modeling mesh dynamics Waczyńska et al. (2024); Gao et al. (2024). To further improve efficiency and generalization, we introduce a sparse voxel-based motion representation that both reduces the complexity of dense vertex deformations and mitigates the synthetic-to-real gap by converting irregular mesh sampling into a unified voxel structure. Moreover, inspired by previous work Li et al. (2024), we further model the motion of each voxel as a spectrum, which can model a long-term mesh motion using a few frequency components, further reducing the complexity for long-term motion generation. At last, we can treat the predicted 3D spectrum as 3D modal bases Li et al. (2024) for modal analysis and approximate 3D interaction by a summation of base motions. By doing so, it reduces the interaction simulation to about $18ms$, making it significantly faster than MPM-based simulation and enabling real-time interaction.
We evaluate our method through comparative experiments on various real-world scenes, showing that our approach produces more natural animations of trees swaying in the wind and dynamically interacting. We summarize our contributions as follows:
  • We introduce DynamicTree, a novel framework for long-term motion generation of real-world trees, enabling realistic swaying animations.
  • We propose a novel sparse voxel spectrum motion representation for efficient and long-term 4D generation. With the generated 3D spectrums, we can further perform fast simulation of interactive dynamics under external forces.
  • To facilitate the generation of complex 3D tree motion, we contribute 4DTree, a large-scale synthetic 4D tree dataset containing 8,786 animated tree meshes, each with 100-frame motion sequences.
@@ 2.1 Tree Animation
Traditional methods for realistic tree animation typically involve creating a physical tree model and simulating its dynamics. For instance, Quigley et al. (2017) represents trees as collections of articulated rigid bodies connected by rotational springs, allowing for flexible yet physically plausible deformations. In contrast, Akagi (2012) introduces a particle-based approach where a link structure between particles replaces the traditional 3D tree model, enabling efficient computation of interaction forces and dynamic responses. Windy-Tree Pirk et al. (2014) incorporates a growth model along with Navier–Stokes equations and the Smoothed Particle Hydrodynamics method. Furthermore, Zhao & Barbič (2013) introduces a semi-automatic pipeline for interactive wind and drag simulations, which, however, requires heavy manual effort to make scanned trees simulation-ready and extensive artistic refinement of materials and lighting during rendering. In summary, although these prior methods can generate realistic dynamics, they remain limited to synthetic physical tree models or suffer from suboptimal rendering, leading to a noticeable gap in visual realism compared with 3DGS-based tree animations.
@@
Recently, 4D content generation has gained growing attention in generative AI. These methods typically construct a static 3D model and then optimize its motion over time. Methods such as MAV3D Singer et al. (2023), Dream-in-4D Zheng et al. (2024), and CT4D Wu et al. (2024) use SDS to generate 3D models from text input, followed by video SDS to animate them. Comp4D leverages large language models (LLMs) Achiam et al. (2023) for motion generation, while 4Dynamic Yuan et al. (2024) bypasses optimization by using generated videos directly. Beyond text-conditioned, Animate124 Zhao et al. (2023) and DreamGaussian4D Ren et al. (2023) combine image-to-3D and video diffusion priors to optimize 4D models. EG4D Sun et al. (2024) and 4DGen Yin et al. (2023) first generate a dynamic video from the input image, and then use multi-view diffusion models Voleti et al. (2024); Liu et al. (2023) to generate multi-view sequences for optimizing the 4D representation. Despite recent progress, these methods rely on 2D motion priors from VDMs due to the lack of real 3D motion data. These limitations often result in inferior temporal and spatial coherence, causing noticeable artifacts in the optimized 3DGS results. Moreover, these approaches depend on scene-specific optimization, entailing substantial computational overhead.
@@ 2.3 Physics-based 3DGS Simulation
Physics-based dynamic generation methods use the differentiable MPM simulation framework to optimize the dynamics of 3DGS. The pioneering work PhysGaussian Xie et al. (2024) employs a customized MPM formulation that bridges Newtonian dynamics and 3D Gaussian kernels, enabling the simulation of various material behaviors. To reduce manual parameter tuning, recent works combine MPM with VDMs or LLMs to estimate physical properties. For instance, PhysDreamer Zhang et al. (2024) and Dreamphysics Huang et al. (2024) integrate motion priors from VDMs with MPM, enabling the learning of dynamic properties such as Young’s modulus and Poisson’s ratio. PhysFlow Comas et al. (2024) initializes parameters via GPT-4 Achiam et al. (2023) and further optimizes them using optical flow guidance. However, as precisely setting individual parameters for each part is challenging, these methods often assume uniform material properties across the entire object to simplify optimization. This facilitates global motion coherence but suppresses local deformation details and reduces visual realism in tree animation. Additionally, the high computational cost of MPM-based simulation limits its use in real-time applications.
@@ 2.4 Spectrum-based Motion Representation
Quasi-periodic motions of plants and trees are well-suited for spectrum-based modeling. Prior work shows they can be modeled as a superposition of a few harmonic oscillators at different frequencies Chuang et al. (2005); Davis (2016); Diener et al. (2009). Generative-Dynamics Li et al. (2024) leverages this property by reconstructing long videos from a few generated frequency components. Moreover, Abe et al. Davis et al. (2015) demonstrate that spectral volumes can serve as image-space modal bases for plausible interactive simulation via modal analysis. Building on this idea, Generative-Dynamics and ModalNeRF Petitjean et al. (2023) apply similar principles to image-space and implicit NeRF representation, achieving interactive dynamic simulations. Inspired by these works, we propose a sparse voxel spectrum representation that enables efficient long-term motion generation and interactive simulation for 3D trees.
@@ 3.1 Task Formulation
Given multi-view images of a static tree, our goal is to generate a 4D model as a deformed 3DGS sequence $\mathcal{G}=\{G_{t}\ |G_{t}=\{x_{i}^{t},r_{i}^{t},s_{i}^{t},\sigma_{i}^{t},c_{i}^{t}\}_{i=1}^{H}\}_{t=0}^{T}$, where $x_{i}^{t},r_{i}^{t},s_{i}^{t},\sigma_{i}^{t}$ and $c_{i}^{t}$ denote the position, rotation, scale, opacity, and color of the $i$-th Gaussian primitive at frame $t$, respectively. This requires predicting temporal deformations of the static 3DGS: $\mathcal{D}_{g}=\{D_{g}^{t}|D_{g}^{t}=(\Delta x_{i}^{t}\in\mathbb{R}^{3},\Delta r_{i}^{t}\in\mathbb{R}^{4},\Delta s_{i}^{t}\in\mathbb{R}^{3})\}_{i=1,t=1}^{H,T}$. Previous methods Yin et al. (2023); Comas et al. (2024) typically rely on optimization-based strategies to solve this problem, which are computationally expensive. We instead formulate this task as a conditional generation problem. To handle large-scale primitives efficiently, we propose a two-stage pipeline, named DynamicTree, as shown in Fig. 2. First, we introduce the sparse voxel spectrum3.2) representation to efficiently represent the motion. Then, we extract voxel grid conditions (§3.3) and employ a sparse voxel diffusion module (§3.4) to generate mesh motion. Subsequently, a two-stage optimization strategy is proposed in §3.5 to refine performance. Finally, we bind 3DGS on the animated mesh surface (§3.6) to compute $\mathcal{D}_{g}$.
@@ 3.2 Sparse Voxel Spectrum
Prior works Wang et al. (2024); Lei et al. (2024) exploit the spatial sparsity of 3D motion using compact bases (e.g., Wang et al. (2024) drives 40k Gaussians with only 20 motion bases). In our case, tree-like motions also exhibit such sparsity. For example, vertices within the same leaf or local branch tend to show similar motion patterns. However, relying solely on sparse motion bases like  Wang et al. (2024) will make it difficult to model fine-grained details due to the complexity of tree motion. Therefore, we propose representing tree motion using sparse voxels, where all vertices in a voxel share the same displacement. With this, to predict dense mesh motion, we only need to predict sparse voxel motion $\mathcal{D}_{v}=\{D_{v}^{t}\in\mathbb{R}^{3\times n}|t=1,...,T\}$, where n is typically an order of magnitude smaller than $N$.
To ensure temporal consistency over long sequences, instead of autoregressive or time-conditioned generation Blattmann et al. (2023); Bertiche et al. (2023); Li et al. (2022), we draw inspiration from Generative-Dynamics Li et al. (2024), which models motion via low-frequency components of spectral volumes Davis et al. (2015). Inspired by this, we introduce the sparse voxel spectrums to represent 3D motion. Specifically, for sparse voxel motion $\mathcal{D}_{v}\in\mathbb{R}^{3\times n\times T}$, we apply the Fast Fourier Transform (FFT) along the temporal dimension, resulting in a complex-valued frequency-domain representation $\hat{\mathcal{D}}_{v}\in\mathbb{C}^{3\times n\times T}$, where each spatial displacement is decomposed into its corresponding components. Tree-like quasi-periodic motions are predominantly captured by the first $K$ frequency components. Thus, a compact representation $\hat{\mathcal{D}}_{v}^{(K)}\in\mathbb{C}^{3\times n\times K}$ is sufficient for nearly lossless reconstruction of the full spectrum $\hat{\mathcal{D}}_{v}\in\mathbb{C}^{3\times n\times T}$, with $K=16$ following Li et al. (2024). Then, to facilitate the generation of these top-$K$ frequency components for the sparse voxel motion, we introduce the sparse voxel spectrum representation $\mathcal{S}=\{s_{i}\in\mathbb{R}^{6\times n}|i=1,...,K\}$, where the dimension of size 6 corresponds to the real and imaginary parts of the $x$, $y$, and $z$ dimensions. Given this representation, we can reconstruct the mesh motion through the following operation:

### v2

@@ Abstract
Generating dynamic and interactive 3D trees has wide applications in virtual reality, games, and world simulation. However, existing methods still face various challenges in generating structurally consistent and realistic 4D motion for complex real trees. In this paper, we propose DynamicTree, the first framework that can generate long-term, interactive 3D motion for 3DGS reconstructions of real trees. Unlike prior optimization-based methods, our approach generates dynamics in a fast feed-forward manner. The key success of our approach is the use of a compact sparse voxel spectrum to represent the tree movement. Given a 3D tree from Gaussian Splatting reconstruction, our pipeline first generates mesh motion using the sparse voxel spectrum and then binds Gaussians to deform the mesh. Additionally, the proposed sparse voxel spectrum can also serve as a basis for fast modal analysis under external forces, allowing real-time interactive responses. To train our model, we also introduce 4DTree, the first large-scale synthetic 4D tree dataset containing 8,786 animated tree meshes with 100-frame motion sequences. Extensive experiments demonstrate that our method achieves realistic and responsive tree animations, significantly outperforming existing approaches in both visual quality and computational efficiency. Intuitive video results are best viewed on our project webpage https://dynamictree-dev.github.io/DynamicTree.github.io/.
Figure 1: DynamicTree achieves structurally consistent and realistic long-term animation and dynamic interaction on real-world 3DGS trees. It first generates mesh motion via a compact sparse voxel spectrum representation and then deforms the surface-bound Gaussian primitives. We visualize the slice of the generated motion at the orange scanline along the time dimension.
@@ 1 Introduction
With recent advances of neural radiance fields (NeRF) [28] and 3D Gaussian Splatting (3DGS) [20], high-quality reconstruction and real-time rendering become feasible. Driven by these, there is high demand to make static reconstructions interactable, for immersive experiences like 3D games, movies, and virtual reality [19, 11, 36]. As a vital component of natural landscapes, tree animation can significantly enrich immersive digital experiences. For example, when viewing a reconstructed backyard on a VR headset, if trees can sway gently in the wind or respond to dragging interaction, it would significantly enhance immersion.
@@
However, animating a realistic 3D tree remains challenging. Traditional tree animation methods [32, 31, 30] typically construct structurally approximated physical tree models and then perform simulation. Although such approaches can simulate realistic motion dynamics, they are either confined to synthetic tree models or require labor-intensive mesh refinements for high-quality rendering. Consequently, they struggle to reproduce the immersive realism and structural complexity of real-world trees, unlike animations based on reconstructed 3DGS representations.
@@
To animate 3DGS trees, existing methods can be broadly categorized into 4D generation and physics-based simulation. 4D generation approaches [3, 38, 24, 56, 50, 27, 52] typically use 2D motion priors from video diffusion models (VDMs) to optimize 4D representations. However, such approaches tend to generate simple motions with temporal and spatial inconsistencies, making realistic tree animation challenging and demanding heavy per-scene optimization. Another category, physics-based approaches, such as PhysGaussian [48], couple 3DGS with physical simulation engines like the Material Point Method (MPM) [18], achieving better 3D consistency. However, they rely on simplified constitutive models and approximate boundary conditions, making fine-grained, realistic branch- and leaf-level motion difficult. These solvers are also computationally expensive, making real-time applications impractical.
@@
In this work, we introduce DynamicTree, a novel 4D generative framework for animating 3D Gaussian Splatting (3DGS) trees with real-time interactivity. Unlike VDM-based 4D generation methods, DynamicTree directly learns tree motion priors in 3D space, avoiding geometric inconsistencies and capturing fine-grained, physically plausible dynamics. Compared with computationally expensive physics-based optimization, our generative model synthesizes fine-grained 3D tree motion in a fast feed-forward manner, achieving more than a hundred times acceleration. To train the model, we construct the first large-scale 4D tree motion dataset containing 8,7868{,}786 animated tree meshes with 100-frame sequences, generated via hierarchical branching simulation [45].
@@
Our framework addresses these challenges with a two-stage pipeline. We first generate mesh motion and then bind Gaussians to the deforming mesh, allowing full 3DGS deformation while only modeling mesh dynamics [43, 12]. To further improve efficiency and generalization, we introduce a sparse voxel-based motion representation that both reduces the complexity of dense vertex deformations and mitigates the synthetic-to-real gap by converting irregular mesh sampling into a unified voxel structure. Moreover, inspired by previous work [23], we further model the motion of each voxel as a spectrum, representing long-term mesh dynamics with only a few frequency components, further reducing temporal complexity. At last, we can treat the predicted 3D spectrum as 3D modal bases [23] for modal analysis and approximate 3D interactions as a linear combination of base motions. By doing so, it reduces the interaction simulation to about $18ms$, making it significantly faster than MPM-based simulation and enabling real-time interaction.
We evaluate our method through comparative experiments on various real-world trees, showing that our approach produces more natural animations of trees swaying in the wind and dynamically interacting. We summarize our contributions as follows:
  • We introduce DynamicTree, a novel framework for 3D motion generation of real-world trees, enabling physically plausible and structurally consistent animations.
  • We propose a novel sparse voxel spectrum representation for efficient long-term 4D tree motion generation and fast interactive simulation under external forces.
  • To support complex 3D tree motion generation, we contribute 4DTree, a large-scale synthetic 4D tree dataset containing 8,786 animated tree meshes, each with 100-frame motion sequences.
@@ 2.1 Tree Animation
Traditional methods for tree animation typically rely on constructing physical tree models and simulating their dynamics. Early works represent trees as articulated rigid bodies or particle-based systems connected by springs or links [32, 2], enabling physically plausible yet simplified deformations. More advanced approaches, such as Windy-Tree [30], couple growth models with fluid dynamics to simulate wind effects, while others introduce semi-automatic pipelines for interactive wind and drag simulations [55]. However, these methods often depend on structurally approximated synthetic physical trees and require extensive manual preparation to make them simulation-ready. Consequently, they struggle to faithfully reproduce the intricate geometry and motion of real-world trees and frequently exhibit suboptimal rendering quality compared with real-world 3DGS-based tree animations.
@@
Recently, 4D content generation has gained growing attention in generative AI. These methods typically optimize 3D motion from 2D diffusion priors. Works such as MAV3D [38], Dream-in-4D [56], and CT4D [47] use SDS-based pipelines to generate and animate 3D content, while Comp4D [1] and 4Dynamic [53] leverage language or video priors for motion synthesis. DreamGaussian4D [34], EG4D [41], 4DGen [52], and SV4D [49, 51] further integrate image-to-3D, video, or multi-view diffusion models [42, 25] for 4D generation. Despite recent advances, these methods depend on 2D motion priors from video diffusion models, leading to temporal and spatial inconsistencies and noticeable artifacts, especially when generating the complex motions of trees. In addition, their reliance on scene-specific optimization incurs substantial computational cost. In contrast, our method learns a 3D diffusion prior from 4D tree data to predict consistent motion in a fast feed-forward manner.
@@ 2.3 Physics-based 3DGS Simulation
Physics-based dynamic generation methods use the differentiable MPM simulation framework to optimize 3DGS dynamics. The pioneering PhysGaussian [48] employs a customized MPM formulation that bridges Newtonian dynamics and 3D Gaussian kernels, enabling the simulation of material behaviors. To reduce manual parameter tuning, recent works combine MPM with VDMs or LLMs to estimate physical properties. For instance, PhysDreamer [54] and Dreamphysics [16], for example, integrate VDM motion priors with MPM to learn dynamic properties such as Young’s modulus and Poisson’s ratio. PhysFlow [26] initializes parameters via GPT-4 [1] and further optimizes them using optical flow guidance. However, as precisely setting individual parameters for each part is challenging, these methods often assume uniform material properties across the entire object to simplify optimization. This facilitates global motion coherence but suppresses local deformation details and reduces visual realism in tree animation. Additionally, the high computational cost of MPM-based simulation limits its use in real-time applications.
@@ 2.4 Spectrum-based Motion Representation
Quasi-periodic motions of plants and trees are well-suited for spectrum-based modeling. Prior work shows they can be modeled as a superposition of a few harmonic oscillators at different frequencies [7, 9, 10]. Generative-Dynamics [23] leverages this property by reconstructing long videos from a few generated frequency components. Moreover, Abe et al. [8] demonstrate that spectral volumes can serve as image-space modal bases for plausible interactive simulation via modal analysis. Building on this idea, Generative-Dynamics and ModalNeRF [29] apply similar principles to image-space and implicit NeRF representation, achieving interactive dynamic simulations. Inspired by these works, we propose a sparse voxel spectrum representation that enables efficient long-term motion generation and interactive simulation for 3D trees.
@@ 3.1 Task Formulation
Given multi-view images of a static tree, our goal is to generate a 4D model as a deformed 3DGS sequence $\mathcal{G}=\{G_{t}\ |G_{t}=\{x_{i}^{t},r_{i}^{t},s_{i}^{t},\sigma_{i}^{t},c_{i}^{t}\}_{i=1}^{H}\}_{t=0}^{T}$, where $x_{i}^{t},r_{i}^{t},s_{i}^{t},\sigma_{i}^{t}$ and $c_{i}^{t}$ denote the position, rotation, scale, opacity, and color of the $i$-th Gaussian primitive at frame $t$, respectively. This requires predicting temporal deformations of the static 3DGS: $\mathcal{D}_{g}=\{D_{g}^{t}|D_{g}^{t}=(\Delta x_{i}^{t}\in\mathbb{R}^{3},\Delta r_{i}^{t}\in\mathbb{R}^{4},\Delta s_{i}^{t}\in\mathbb{R}^{3})\}_{i=1,t=1}^{H,T}$. Previous methods [52, 26] typically rely on optimization-based strategies to solve this problem, which are computationally expensive. We instead formulate this task as a conditional generation problem. To handle large-scale primitives efficiently, we propose a two-stage pipeline, named DynamicTree, as shown in Fig. 2. First, we introduce the sparse voxel spectrum3.2) representation to efficiently represent the motion. Then, we extract voxel grid conditions (§3.3) and employ a sparse voxel diffusion module (§3.4) to generate mesh motion. Subsequently, a two-stage optimization strategy is proposed in §3.5 to refine performance. Finally, we bind 3DGS on the animated mesh surface (§3.6) to compute $\mathcal{D}_{g}$.
@@ 3.2 Sparse Voxel Spectrum
Prior works [44, 21] exploit the spatial sparsity of 3D motion using compact bases (e.g., [44] drives 40k Gaussians with only 20 motion bases). In our case, tree-like motions also exhibit such sparsity. For example, vertices within the same leaf or local branch tend to show similar motion patterns. However, relying solely on sparse motion bases like  [44] will make it difficult to model fine-grained details due to the complexity of tree motion. Therefore, we propose representing tree motion using sparse voxels, where all vertices in a voxel share the same displacement. With this, to predict dense mesh motion, we only need to predict sparse voxel motion $\mathcal{D}_{v}=\{D_{v}^{t}\in\mathbb{R}^{3\times n}|t=1,...,T\}$, where n is typically an order of magnitude smaller than $N$.
To ensure temporal consistency over long sequences, instead of autoregressive or time-conditioned generation [5, 4, 22], we draw inspiration from Generative-Dynamics [23], which models motion via low-frequency components of spectral volumes [8]. Inspired by this, we introduce the sparse voxel spectrums to represent 3D motion. Specifically, for sparse voxel motion $\mathcal{D}_{v}\in\mathbb{R}^{3\times n\times T}$, we apply the Fast Fourier Transform (FFT) along the temporal dimension, resulting in a complex-valued frequency-domain representation $\hat{\mathcal{D}}_{v}\in\mathbb{C}^{3\times n\times T}$, where each spatial displacement is decomposed into its corresponding components. Tree-like quasi-periodic motions are predominantly captured by the first $K$ frequency components. Thus, a compact representation $\hat{\mathcal{D}}_{v}^{(K)}\in\mathbb{C}^{3\times n\times K}$ is sufficient for nearly lossless reconstruction of the full spectrum $\hat{\mathcal{D}}_{v}\in\mathbb{C}^{3\times n\times T}$, with $K=16$ following [23]. Then, to facilitate the generation of these top-$K$ frequency components for the sparse voxel motion, we introduce the sparse voxel spectrum representation $\mathcal{S}=\{s_{i}\in\mathbb{R}^{6\times n}|i=1,...,K\}$, where the dimension of size 6 corresponds to the real and imaginary parts of the $x$, $y$, and $z$ dimensions. Given this representation, we can reconstruct the mesh motion through the following operation: